If an ODE is given as

(x2+x+1)dx+(y2-y+3)dy=0 with y(0)=1.

Find the numerical solution with h=0.1.



1
Expert's answer
2021-03-18T15:43:22-0400

Solution

For the first-order differential equation with the initial value

y’(x) = F(x,y y(x)),  y(x0) = y0  

according to the Euler method

yn+1 = yn + h*F(xn, yn)

where h – step, xn = h*n, yn = y(xn), n = 0,1,2…N

In this case F(x,y) = -(x2+x+1)/(y2-y+3) , h = 0.1 , x0 = 0 , y0 = 1.

Calculations gives such results:

n        xn           yn

0        0.0         1.000

1        0.1         0.967

2        0.2         0.929

3        0.3         0.887

4        0.4         0.839

5        0.5         0.785

6        0.6         0.723

7        0.7         0.653

8        0.8         0.574

9        0.9         0.485

10      1.0         0.387